# Configuration Interaction Guided Sampling with Interpretable Restricted Boltzmann Machine

### v1

@@ Abstract
We propose a data-driven approach using a Restricted Boltzmann Machine (RBM) to solve the Schrödinger equation in configuration space. Traditional Configuration Interaction (CI) methods, while powerful, are computationally expensive due to the large number of determinants required. Our approach leverages RBMs to efficiently identify and sample the most significant determinants, accelerating convergence and reducing computational cost. This method achieves up to 99.99% of the correlation energy even by four orders of magnitude less determinants compared to full CI calculations and up to two orders of magnitude less than previous state of the art works. Additionally, our study demonstrate that the RBM can learn the underlying quantum properties, providing more detail insights than other methods . This innovative data-driven approach offers a promising tool for quantum chemistry, enhancing both efficiency and understanding of complex systems.
keywords:
@@
Full Configuration Interaction (FCI); Restricted Boltzmann Machine (RBM); Data-Driven Quantum Chemistry.
journal: Computer Physics Communications
1 Introduction
Understanding the quantum properties of molecules and materials is crucial in various scientific and technological fields. The Schrödinger equation serves as the fundamental base for this understanding, but its exact solution is often intractable. The Configuration Interaction (CI) method offers a powerful approximation, but it suffers from the fact that the vast number of determinants are required for realistic systems. This complexity limits the applicability of traditional CI methods to many important molecules.
This work proposes a novel approach that harnesses the power of artificial intelligence, specifically Restricted Boltzmann Machines (RBMs), to address the limitations of CI. RBMs are probabilistic neural networks that are experts at efficiently sampling prohibitive spaces to find the current distribution. In recent years have been used in the field of Computational Biology Medicine and Chemistry in pattern extraction and learning complex distributions [1, 2, 3].
Although new generative models have been developed in artificial intelligence such as Generative Adversarial Networks (GAN), Variational Autoencoders(VAE) or Deep learners, RBMs have become popular because they only have one visible and one hidden layer, resulting in simple and potentially interpretable learned features.
Our strategy uses the fact that, like Principal Component Analysis (PCA), an RBM can learn the principal components that maximize the variance in a dataset. This allows us to obtain the main characteristics. However, unlike PCA, which has linear transformations of the data vectors, RBMS are non-linear in nature and have the ability to learn more complicated relationships between visible and hidden units. In this work, we use an RBM to identify the most relevant determinants in the configuration space, accelerating convergence in solving the Schrödinger equation compared to random sampling or traditional CI methods.
The determinants, are fundamental to CI methods, such determinants are governed by Slater’s rules [4]. These rules dictate that only determinants differing by one or two permutations have non-zero Hamiltonian matrix elements. This implies an inherent relationship among the determinants, suggesting that those making the most significant contributions to the wavefunction share common features. Our RBM-based approach exploits this relationship by learning the underlying distribution of these determinants that contribute more to the solution of the Schrödinger equation.
It is noteworthy that the CI method can be approximated using various levels of theory, such as configuration interaction with single and double excitations (CISD) and considering triple, and quadruple excitations too (CISDTQ). CISD, for instance, can recover approximately $95\%$ of the correlation energy (in small molecules at equilibrium geometries) [5, 6], while CISDTQ can achieve up to $99.8\%$ of the full CI energy [7]. These approximations highlight the importance of employing higher-order methods like Full Configuration Interaction (FCI) to obtain more accurate results.
Traditional approaches to CI, such as Quantum Monte Carlo or stochastic search methods [8], can be computationally inefficient. In contrast, RBMs, while incorporating a stochastic process at their core (Gibbs Sampling or Contrastive Divergence), ultimately guides the sampling process based on the learned data distribution. This guided approach enhances the efficiency and effectiveness of the RBM-based CI method.
Previous works has explored the use of RBMs to solve the CI approximation [9, 10]. In this study, we present a significant improvement to the state of the art by introducing an innovative and differentiated model training methodology.
Specifically, our approach utilizes RBM as generative model with a innovative guided model strategy that focuses on iteratively selecting determinants with higher probabilities, thereby facilitating faster and more precise solution convergence.
A key result of this study is the ability to reach similar or even better results as current state of the art methods, with up to two orders of magnitude less determinants. Such improvement is obtained by implementing a taboo list that allows for retaining only the determinants that contribute the most to the final wavefunction in each iteration, and using that information for subsequent training when generating new determinants while searching the Hilbert space.
This strategy helps us maintain the time complexity and prevent rapid growth in computational costs, reducing overall computation time.
Furthermore, we present, a novel guided learning strategy for the RBM model, and demonstrate that such model is explainable. The insights learned from the RBM model provide a deeper understanding of the underlying patterns that make up the solution. This investigation revealed that the model appears to discern a distinctive pattern across various molecules, which remains consistent within the same molecule using different molecular basis. This pattern seems linked to the radial distribution of probability of specific molecular bonds and the energies of energy orbitals. This suggests that the model has uncovered a previously unrecognized relationship among determinants. Such interpretability not only enhances current applications but also lays the foundation for future research in applying artificial intelligence techniques to quantum chemistry.
In summary, this study represents a significant advance in the field of computational quantum chemistry by introducing an improved CI solution strategy, supported by a guided RBM learning model, accompanied by a practical implementation and a detailed analysis of the results.
The paper begins with Section 2, which provides a theoretical background on solving the Schrödinger equation for complex systems (2.1). It also covers the fundamentals of the Configuration Interaction method and its limitations (2.2), along with an overview of Restricted Boltzmann Machines (2.3).
In Subsection (2.4), we introduce our proposed model that utilizes an RBM as a generative model, highlighting key features and innovations aimed at approximating FCI. We then present the results and discussions (3), followed by the conclusions in Section (4).
2 Material and methods
2.1 Many-body Hamiltonian
In quantum mechanics, solving the time-independent Schrödinger equation (Eq. 1) is the cornerstone for predicting the behavior of a system:
$H\psi=E\psi.$ (1)
Here, $H$ represents the system’s Hamiltonian, $\psi$ is the wave function, and $E$ is the corresponding energy eigenvalue. Solving this eigenvalue equation provides the complete energy spectrum and wave functions, enabling the calculation of all properties and predictions for the system.
However, Eq. 1 presents a significant challenge when dealing with many-body problems. Consider a solid-state system with a vast number of microscopic constituents (atoms) and constituent particles (electrons). An accurate description requires a Hamiltonian that incorporates the kinetic energies of both electrons and nuclei, along with all interaction terms: electron-electron, electron-nuclei, and nuclei-nuclei. This leads to a more comprehensive form of the Schrödinger eigenvalue equation given by Eq. 2:
$\left[\frac{h^{2}}{2m}\sum_{i=1}^{N}\nabla_{i}^{2}+\sum_{i=1}^{N}V(r_{i})+\right.\left.\sum_{i=1}^{N}\sum_{j<1}U(r_{i},r_{j})\right]\psi=E\psi.$ (2)
Here, $m$ represents the electron mass. The first term describes the kinetic energy of each electron due to their individual momenta. The second term, $V(r_{i})$, represents the potential energy of each electron interacting with the collection of atomic nuclei, often referred to as the external potential. This term can be further expanded as:
$\sum_{i=1}^{N}V(r_{i})=\sum_{i=1}^{N}\sum_{j=1}^{N_{atoms}}\frac{-Z_{j}}{|r_{i}-R_{j}|},$ (3)
where $Z_{j}$ is the atomic number (number of protons) of the $j$-th nucleus, and $r_{i}$ and $R_{A}$ are the position vectors of the $i$-th electron and $j$-th nucleus, respectively. The negative sign indicates the attractive nature of the interaction between negatively charged electrons and positively charged nuclei.
Considering the Born-Oppenheimer approximation, the nuclei are considered stationary due to their much larger mass compared to electrons. Therefore, they exert a fixed Coulombic potential on the electrons [11].
The third term in the Hamiltonian, $U(r_{i},r_{j}$), represents the potential energy arising from the interaction between different electrons due to their Coulombic repulsion. This term is expressed as:
$\sum_{i=1}^{N}\sum_{j<1}U(r_{i},r_{j})=\sum_{i=1}^{N}\sum_{j<1}\frac{e^{2}}{2|r_{i}-r_{j}|}.$ (4)
In Eq. 2, $\psi$ represents the electronic wave function, a function that depends on the spatial coordinates of all $N$ electrons: $\psi=\psi(r_{1},...r_{n})$. The solutions of the Schrödinger equation with this Hamiltonian provide the corresponding energy levels ($E$) of the electronic system.
The complexity arises due to the high dimensionality of the electronic wave function. For instance, a single CO2 molecule requires a 66-dimensional wave function (3 spatial coordinates for each of the 22 electrons). For a larger system, like a nanocluster with 100 Pt atoms, the wave function becomes a staggering 23,000-dimensional object. This immense complexity makes it computationally intractable to solve the Schrödinger equation exactly for most many-body systems [12].
2.1.1 Challenges and Approximation Methods:
Due to the high dimensionality of the many-body problem, various approximation methods are employed to solve the Schrödinger equation. Two of the most well-established techniques are wave function-based methods and density-based methods. Wave function-based methods include perturbative and variational approaches, such as the Configuration Interaction (CI) method, which will be the primary focus of the following section. CI explicitly considers a linear combination of multiple electronic configurations to improve the accuracy of the wave function.
On the other hand, Density Functional Theory (DFT) is a popular density-based method known for its computational efficiency. However, DFT’s application can still suffer from large pervasive errors that cause qualitative failures in predicted properties. For instance, DFT often underestimates the barriers of chemical reactions, the band gaps of materials, the energies of dissociating molecular ions, and charge transfer excitation energies. Additionally, it tends to overestimate the binding energies of charge transfer complexes and the response to an electric field in molecules and materials [13].
In contrast, CI, particularly Full Configuration Interaction (FCI), offers a rigorous wave function-based approach that does not rely on approximations of the exchange-correlation functional. This makes CI a preferred choice in cases where the highest accuracy is required [14], such as in the study of bond-breaking processes, excited states, and systems with near-degenerate electronic states. Despite its higher computational cost, CI provides a more detailed and accurate description of electron correlation effects.
We will delve deeper into the details and applications of CI, specifically FCI, in the next section.
2.2 Configuration Interaction
There are well-established methods for approximating the wave function in order to determine the ground state of electronic systems. One prominent method is the Hartree-Fock method, which simplifies the problem by assuming that each electron moves independently in a central potential created by the other electrons. This eliminates the need to consider the $r_{i}$ term in the many-body Hamiltonian equation. However, the accuracy of the ground state energies obtained using this method is limited due to the use of an average potential to account for electron-electron interactions [6].
To overcome this limitation, Configuration Interaction (CI) methods are employed. These methods acknowledge that electrons interact and repel each other, leading to correlated motion within atoms. Instead of being restricted to a single orbital configuration represented by a Slater determinant, CI methods construct properly antisymmetrized product wave functions using multiple configurations [15]. By distributing electrons among different orbitals, spatial correlations in electron motion can be taken into account. CI utilizes a variational wave function [16] mathematically expressed as shown in Eq. 5:
$\Psi=\sum_{i=0}^{I}c_{i}\Phi_{i}=c_{0}\Phi_{0}+c_{1}\Phi_{1}+\ldots$ (5)
In this expansion of the total wave function ($\Psi$), it is expressed as a linear combination of diferent electron configurations ($\Phi_{i}$).The bases used in the expansion are usually excited Slater determinants [17] . The expansion coefficients ($c_{i}$) determine the contribution of each configuration to the total wave function.
The Hamiltonian matrix is constructed so that $H_{ij}=\langle\Phi_{i}|\hat{H}|\Phi_{j}\rangle$ for $i,j=0,1,2,\ldots,I$. Using Slater’s rules allows us to express the matrix elements $H_{ij}$ in terms of one- and two-electron integrals. The one-electron integrals are written as [15]:
@@ 2.2 Configuration Interaction
Here, $\phi_{i}$ and $\phi_{j}$ are spin-orbitals, which are functions of both spatial and spin coordinates of a single electron. The Hamiltonian is then diagonalized to obtain the coefficients $c_{i}$ in the expansion of the wave function.
When the expansion in Eq. 5 includes all possible electron configurations with the appropriate symmetry, it constitutes a full configuration interaction (FCI) procedure that exactly solves the electronic Schrödinger equation within the space spanned by the one-particle basis set.
If only determinants of single and double excitations are considered, the method is known as Configuration Interaction Singles Doubles (CISD) [18]. All combinations, or every possible determinant is referred to as Full Configuration Interaction (FCI), providing the most accurate approximation to the ground state within the specific basis set. However, the FCI approach as a complete solution, suffers from poor scalability as the number of determinants grows exponentially with $\binom{2K}{N}$ [18], where $2K$ is the number of spin-orbitals and $N$ is the number of particles [10].
The total number of Slater determinants, $N_{det}$, is calculated as follows (assuming a closed-shell singlet state):
$N_{det}=\left(\frac{N_{o}!}{(N_{o}-N_{e})!N_{e}!}\right)^{2}\approx\binom{N_{o}}{N_{e}/2}^{2}$ (8)
where $N_{o}$ represents the number of molecular orbitals (MOs) and $N_{e}=N_{\alpha}=N_{\beta}$ represents the number of electrons with $\alpha$-spin, which is half of the total number of electrons [19]. This formula highlights the rapid growth in the number of determinants as the size of the system increases, contributing to the computational challenges of the FCI approach.
The following Table 1 shows the number of determinants required for simple molecules, demonstrating that a significant number of determinants are needed even for these simple systems.

### v2

@@ Abstract
We propose a data-driven approach using a Restricted Boltzmann Machine (RBM) to solve the Schrödinger equation in configuration space. Traditional Configuration Interaction (CI) methods construct the wavefunction as a linear combination of Slater determinants, but this becomes computationally expensive due to the factorial growth in the number of configurations. Our approach extends the use of a generative model such as the RBM by incorporating a taboo list strategy to enhance efficiency and convergence. The RBM is used to efficiently identify and sample the most significant determinants, thus accelerating convergence and substantially reducing computational cost. This method achieves up to 99.99% of the correlation energy while using up to four orders of magnitude fewer determinants compared to full CI calculations and up to two orders of magnitude fewer than previous state of the art methods. Beyond efficiency, our analysis reveals that the RBM learns electron distributions over molecular orbitals by capturing quantum patterns that resemble Radial Distribution Functions (RDFs) linked to molecular bonding. This suggests that the learned pattern is interpretable, highlighting the potential of machine learning for explainable quantum chemistry
@@
Keywords: Full Configuration Interaction (FCI);Machine Learning for electronic structure; Restricted Boltzmann Machine (RBM); Data-Driven Quantum Chemistry; interpretable generative models
I.   Introduction
Understanding the quantum properties of molecules and materials is essential across many scientific and technology fields. While the Schrödinger equation forms the theoretical foundation for this understanding, its exact solution is often intractable for more than a few electrons. Configuration Interaction (CI) (Karazija and Kučas, 2013, Bunge, 2010, Hussein, 2015) methods offer a powerful approximation by expanding the wavefunction in a basis of electronic configurations. However, the number of determinants grows factorially with the number of electrons and molecular orbitals, which limits their practical applicability.
To address this challenge, we propose a novel approach that leverages Restricted Boltzmann Machines (RBMs) to efficiently sample the configuration space, focusing on those determinants with the highest contribution to the solution. RBMs are probabilistic neural networks with a simple two-layer structure that can learn complex, nonlinear distributions. Unlike other generative models such as Generative Adversarial Networks (GANs) or Variational Autoencoders (VAEs), the RBM’s simple structure allows for efficient sampling and makes it easier to interpret the learned features. Our model extends the use of the RBM for wavefunction solutions (Carleo and Troyer, 2017, Herzog et al., 2022), building upon their generative capabilities to enhance sampling efficiency and interpretability.
The key innovation of our approach lies in its iterative, guided model training strategy. At each iteration, determinants with higher contributions (as governed by Slater’s rules) are prioritized and retained, while a taboo list prevents the re-selection of low impact determinants. This not only reduces the number of determinants required by up to two orders of magnitude but also keeps the computational cost manageable. Additionally, the method provides interpretability: The RBM appears to uncover distinctive patterns in the distribution of determinants that correlate with physical properties such as orbital energies and radial probability distributions.
In summary, our study introduces a significant advance in computational quantum chemistry by integrating a guided RBM-based CI strategy with a practical implementation, achieving a reduction in computational cost compared to traditional and state of art CI methods. The remainder of the paper is organized as follows. Section 2 provides the theoretical background on solving the Schrödinger equation and reviews the fundamentals and limitations of CI methods, along with an overview of RBMs. In Subsection III, we detail our proposed model and its key innovations. Section IV presents the results and discussion, and Section V concludes the paper.
II.   Methodology
A.   Theoretical Background
In quantum mechanics, the time-independent Schrödinger equation,
$H\psi=E\psi,$
provides the fundamental framework to predict a system’s properties. Here, $H$ is the Hamiltonian operator, $\psi$ is the many-electron wave function, and $E$ represents the energy eigenvalue. For many-body systems, an accurate Hamiltonian must account for the kinetic energies of electrons and nuclei as well as their mutual interactions. Under the Born–Oppenheimer approximation, the nuclei are treated as fixed, leading to a Hamiltonian of the form:
$\left[\frac{\hbar^{2}}{2m}\sum_{i=1}^{N}\nabla_{i}^{2}+\sum_{i=1}^{N}V(r_{i})+\sum_{i=1}^{N}\sum_{j<i}U(r_{i},r_{j})\right]\psi=E\psi.$
In this equation, the first term represents the kinetic energy of each electron, the second term $V(r_{i})$ describes the Coulomb attraction between electrons and nuclei,
$\sum_{i=1}^{N}V(r_{i})=\sum_{i=1}^{N}\sum_{j=1}^{N_{\text{atoms}}}\frac{-Z_{j}}{|r_{i}-R_{j}|},$
and the third term $U(r_{i},r_{j})$ accounts for electron-electron repulsion:
$\sum_{i=1}^{N}\sum_{j<i}U(r_{i},r_{j})=\sum_{i=1}^{N}\sum_{j<i}\frac{e^{2}}{|r_{i}-r_{j}|}.$
Due to the rapidly increasing dimensionality of the wave function, the calculation of an exact solution of the Schrödinger equation becomes computationally intractable. For instance, a CO2 molecule involves a 66-dimensional wave function, while a nanocluster of 100 Pt atoms may require over 23,000 dimensions (Sholl and Steckel, 2011).
To overcome this “curse of dimensionality,” several approximation methods have been developed. For example, one common strategy is to use wave function-based approaches that improve upon a reference state. This reference is typically provided by the Hartree–Fock (HF) method, which approximates the many electron problem by representing the system with a single Slater determinant and treating electron–electron interactions through a mean field potential. However, HF neglects dynamic electron correlation and often provides only a qualitative description of the system (Marti-Dafcik et al., 2025, Booth et al., 2009, Tubman et al., 2016).
To account for these correlation effects, post Hartree–Fock methods expand the wave function as a linear combination of multiple electronic configurations, also known as configuration state functions (CSFs):
$\Psi=\sum_{i=0}^{I}c_{i}\Phi_{i}=c_{0}\Phi_{0}+c_{1}\Phi_{1}+\ldots,$ (1)
where each CSF $\Phi_{i}$ is weighted by a coefficient $c_{i}$ that indicates its contribution to the total wave function. Importantly, these coefficients are not uniformly distributed; only a small subset of determinants contribute in a significant way. Thus, identifying these key determinants is essential for achieving a good approximation with fewer configurations.
Truncated Configuration Interaction (CI) methods form a hierarchy of approximations based on limiting the excitation rank relative to the Hartree–Fock reference. For example, CISD includes only single and double excitations, while CISDTQ extends this to triple and quadruple excitations. Each successive level captures increasingly subtle electron correlation effects. For example, CISD typically recovers about 95% of the correlation energy in small molecules at equilibrium geometries (Fermann et al., 1994, Blinder, 2019), while CISDTQ can approach 99.8% of the full CI energy (Tschumper and Schaefer, 1998). At the top of this hierarchy lies Full Configuration Interaction (FCI), which includes all possible excitations within a given one-particle basis and provides, in principle, an exact solution to the Schrödinger equation in that space. However, the factorial growth in the number of determinants with system size, makes FCI computationally infeasible for anything beyond the smallest molecules.
Coupled Cluster methods, particularly coupled cluster singles doubles perturbative triples CCSD(T), also rely on a determinant based expansion, but use an exponential ansatz that offers size extensivity and excellent accuracy for ground state properties. CCSD(T) is often regarded as the ”gold standard” in quantum chemistry due to its reliable performance across a wide range of molecular systems (Bartlett and Musiał, 2007). However, it remains computationally expensive, with a scaling of $\mathcal{O}(N^{7})$ with respect to system size, and it may fail in situations involving strong static correlation or near degenerate electronic states, such as bond dissociation or transition metal complexes (Hait et al., 2019, Bulik et al., 2015)
Another notable approach is the Configuration Interaction using a Perturbative Selection made Iteratively (CIPSI). CIPSI iteratively adds determinants to the Hamiltonian by selecting those with the largest second order perturbative contributions, thus efficiently capturing the most important correlation effects without exhaustively exploring the entire configuration space. However, CIPSI relies heavily on perturbation theory approximations, which can lead to inaccuracies when describing strongly correlated systems or when the second-order perturbative corrections become large, thus limiting its reliability and computational efficiency in some challenging cases Dash et al. (2021).
Alternatively, density-based methods such as Density Functional Theory (DFT) (Cohen et al., 2008) offer computational efficiency but may suffer from systematic errors (e.g., underestimating reaction barriers or band gaps (Szabo and Ostlund, 1996b)).
Because the most important contributions to $\Psi_{\text{FCI}}$ are concentrated in a small subset of determinants, it is crucial to develop methods that can selectively sample and include only those determinants. This motivates our approach, which uses artificial intelligence, specifically, a RBM to guide the selection of determinants, thereby reducing computational cost while preserving accuracy. In the following subsection, we provide a brief overview of the theoretical foundations of FCI and the RBM model that supports our method.
B.   Background on FCI and RBM
In electronic structure calculations, solving the Schrödinger equation is often done using well established methods that represent the many electron wavefunction as a linear combination of electronic configurations. While the Hartree Fock method simplifies the problem by assuming that each electron moves independently in a central potential, neglecting the explicit $r_{i}$ term in the many body Hamiltonian and limiting accuracy due to the use of an average potential for electron-electron interactions (Blinder, 2019). Therefore CI methods overcome this limitation by explicitly accounting for electron correlation.Thus, CI constructs properly antisymmetric wave functions by combining multiple electronic configurations, typically represented by excited Slater determinants (Equation 1). Each Slater determinant is an antisymmetric function of spin orbitals, ensuring compliance with the Pauli exclusion principle (David Sherrill and Schaefer, 1999, Szabo and Ostlund, 1996a). In practice, the Hamiltonian matrix is constructed by evaluating its elements using Slater’s rules, which require the computation of one and two-electron integrals:
@@ B.   Background on FCI and RBM
which populate the Hamiltonian elements
$H_{ij}=\langle\Phi_{i}|\hat{H}|\Phi_{j}\rangle,$
that are then diagonalized to obtain the coefficients $c_{i}$ (see Eq. 1) (David Sherrill and Schaefer, 1999, Lestrange et al., 2018).
Although the factorial growth of the determinant space is well known, only a small subset of configurations contributes significantly to the wavefunction, as the coefficients $c_{i}$ are highly not uniformly distributed. Each configuration represents a unique arrangement of electrons among the available molecular orbitals, whose number depends on the chosen basis set. For instance, in a closed shell system, the configurations correspond to the possible distributions of alpha (spin up) and beta (spin down) electrons, always obeying the Pauli exclusion principle (i.e., no two electrons with the same spin can occupy the same orbital). As a result, the full configuration space quickly becomes intractable. For example, the estimated number of determinants is given by
$N_{det}\approx\binom{N_{o}}{N_{e}/2}^{2},$
where $N_{o}$ is the number of molecular orbitals and $N_{e}$ is the total number of electrons. This estimation leads to the values shown in Table 1 for simple molecules.